Spectrum:

If I use distributive property to solve this problem how would I do it? 1(2x+4) -2+2=10

Spectrum:

We just started learning this and I just be sitting in the back of class staring like 0-0

Extrinix:

Ok, so first off do you know what you have to do first?

kittybasil:

Rewording:\[1(2x+4)-2+2=10\]Firstly, you can see that two numbers cancel out:\[1(2x+4)\cancel{\color{red}{-2+2}}=10\]\[\therefore1(2x+4)=10\]Now we apply the distributive property, which is \(a(b+c)=a(b)+a(c)\). Based on that, what should the next step be?

Extrinix:

Annnd I got canceled out... as usual...

kittybasil:

kittybasil wrote:
Rewording:\[1(2x+4)-2+2=10\]Firstly, you can see that two numbers cancel out:\[1(2x+4)\cancel{\color{red}{-2+2}}=10\]\[\therefore1(2x+4)=10\]Now we apply the distributive property, which is \(a(b+c)=a(b)+a(c)\). Based on that, what should the next step be?
Hint: \(b=2x\) I know, usually when you start working with the Distributive property each placeholder variable is a single-digit term, but now we're going with algebraic expressions and variable unknowns.

kittybasil:

extrinix wrote:
Annnd I got canceled out... as usual...
nah, you've got a medal from OP. I'm just here as the tip machine. πŸ˜‚

kittybasil:

I don't want that medal so don't throw it this way! runs away quickly

Spectrum:

So 2x+4=10

Spectrum:

subtract 4 from both sides, 2x=6

kittybasil:

Correct! You're doing great πŸŽ‰

Spectrum:

divide both sides by 2 and x=3

Spectrum:

thenkyou extrinix and basil

Extrinix:

No need to thank me...

kittybasil:

Absolutely thank him. He was here first. Also I go by Kitt haha

kittybasil:

Good luck on your studies! ⭐

kittybasil:

extrinix wrote:
I'm going to swap your name around a bit, Bay.
No pls. I only go by my full user or Kitt. πŸ˜”

Spectrum:

Thenkyou! β€οΈπŸ”’ no need to make this post off-topic. so plez dont❀️❀️

kittybasil:

It's okay. Just clarifying what people should use to refer to me is all. 😊